Bronze which costs $9.10/kg is made by combining copper which costs $8.90/kg with tin which costs $9.50/kg. Find the number of kg of copper and tin required to make 15.3 kg of bronze.
Thank You,
Shubham

Answer by scott8148(6628)

c + t = 15.3

8.9c + 9.5t = 9.1 * 15.3

solve by substitution or elimination